f(x) = -3x^2 + x, x = -5
Answer by nerdybill(7384)

f(x) = -3x^2 + x, x = -5
.
They want to know the value of the function f(x) when x = -5
So, simply substitute the -5 for the 'x' and solve:
f(-5) = -3(-5)^2 + (-5)
f(-5) = -3(25) + (-5)
f(-5) = -75 - 5
f(-5) = -80